OpenStudy (katielong):

A=^3√36V² rearrange so v is subject

OpenStudy (saifoo.khan):

\[A=\sqrt[3]{36v^2}\]THat?

OpenStudy (katielong):

yeh, sorry

OpenStudy (saifoo.khan):

take cube on both sides. \[A^3 = 36 v^2\]

OpenStudy (saifoo.khan):

Multiply 36 on both sides, \[\frac{A^3}{36}=v^2\]

OpenStudy (saifoo.khan):

Now, take a square root on both sides to cancel the v^2 So, the final answer will be: \[v=\sqrt{\frac{A^3}{36}}\]

OpenStudy (katielong):

i got that but brackets around \[A^3/36\] instead, thanks

OpenStudy (saifoo.khan):

Oh, you welcome.